OpenStudy (anonymous):

If the revenue of a shoe company is f(x) = 33x + 16, where x is the number of pairs of shoes sold, and the expenses of the company are g(x) = 4x + 50, where x is also the number of pairs of shoes sold, what is the profit of the company if they sell 105 pairs of shoes? A. $3,481 B. $3,011 C. $3,951 D. $3,111

OpenStudy (anonymous):

B. $3,011 because you want to subtract the expense equation from the revenue equation to find your profit. Substitute 105 as the x in the equations. So for the Revenue f(x)=33(105)+16. The Expense is g(x)=4(105)+50. Follow order of operations for both equations so multiply first then add or subtract. Once you find what the Revenue and Expense equals subtract the expenses from the revenue and you should end up with $3,011 for the answer.

OpenStudy (anonymous):

B. $3,011. because you want to subtract the expense equation from the revenue equation to find your profit. Substitute 105 as the x in the equations. So for the Revenue f(x)=33(150)+16. The Expense is g(x)=4(105)+50. Follow order of operations for both equations so multiply first then add or subtract. Once you find what the Revenue and Expense equals subtract the expenses from the revenue and you should end up with $3,011 for the answer.